Three identical balls with masses of M, 2M, and 3M are fastened to a massless rod of length L as shown. The rotational inertia about the left end of the rod is ?

Please help. Thank you


1
Expert's answer
2022-03-07T07:16:46-0500

We know that

"I=mr^2"

"I=I_1+I_2+I_3"


"I=(3M)\times0^2+2M\times(\frac{L}{2})^2+ML^2"

"I=\frac{3}{2}ML^2"